# Robust Few-shot Transfer Learning for Knowledge Base Question Answering with Unanswerable Questions

### v1

@@ 3.2 Proposed Approach: FUn-FuSIC
In the absence of a clear winner (i.e., a logical form that passes all checks), FUn constructs a candidate set $L$ containing logical forms that are flawed, but not fatally so. Logical forms that fail C1, C2, C4a, or C4b are discarded. Those that fail C3 but pass C4c (have non empty answer but disagree with the question), or fail C4c but pass C3 (have empty answer but are equivalent to the original question) are included in $L$.
@@
Consistency breaks down when all the reasoning paths have low probability, leading to a diverse set of answers with little consensus. For our candidate set $L$ with potentially flawed logical forms, the most popular answer $a^{*}$ may have very few votes. Self consistency for unanswerability (scUn) returns the most popular answer $a^{*}$, only if it receives at least $t$ votes out of $n$. We use $t=\frac{n}{2}$. Otherwise, it returns no consistency.
When there is consistency, scUn outputs the most consistent answer $a^{*}$, empty or non-empty. However, we additionally need to identify the most likely logical form with answer $a^{*}$. When $L(a^{*})\subset L$ has more than one logical form, we use LLM prompting to select the most likely $l^{*}\in L(a^{*})$. Details are in the Appendix.(Sec. A.2)
Adapting FUn for Answerable KBQA:
Though motivated by unanswerability, the core idea of iterative generation of candidate logical forms with feedback from error checks, followed by consistency check is applicable more broadly. To adapt this for answerable KBQA, we make the following simple modifications. (a) Logical forms that fail C4c (empty answer) are not included in the candidate set $L$. (b) Instead of scUn, we fall back on regular self consistency for $L$, so that a consistent answer is always returned. These imply that ($l=l^{*}$, $a=a^{*}$) is the only possible outcome.
@@ 4 Experiments
FuSIC-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
FUn-FuSIC 76.1 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
@@
We now present the experimental evaluation of FUn-FuSIC. First, we aim to address the following research questions for the few-shot transfer setting for KBQA with unanswerability. (R1) How does FUn-FuSIC compare against SoTA models for KBQA with unanswerability, adapted for the few-shot setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) To what extent do the different components of FUn-FuSIC contribute to its performance? In addition, we ask the following research question for the answerable KBQA setting. (R4) How does FUn-FuSIC compare against SoTA few-shot transfer (answerable) KBQA models?
@@ 4.1 Experimental Setup
We create our target test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from the GrailQAbility and GraphQAbility test sets. Additionally, we create the target few shots by selecting 50 answerable and 50 unanswerable questions uniformly at random from the GrailQAbility dev set and the GraphQAbility train set respectively. Of the answerable questions, schema level unanswerable forms 67% in GrailQAbility and 75% in GraphQAbility. The source WebQSP training set has 2858 questions, all of which are answerable.
@@
Table 2: Performance of various models for different categories of unanswerable questions in two datasets.
@@
There are no existing few-shot transfer KBQA models for unanswerability. RetinaQA Faldu et al. (2024) is the SoTA in-domain model for KBQA with unanswerability. It has been shown to significantly outperform various unanswerability adaptations of all other SoTA in-domain KBQA models, such as RnG-KBQA Ye et al. (2022), TIARA Shu et al. (2022), and Pangu Gu et al. (2023), across different categories of unanswerability. We use the available code for RetinaQA. [^2] https://github.com/dair-iitd/RetinaQA We train all components on WebQSP, using as dev set the corresponding target domain dev set. In the absence of unanswerable questions for training, RetinaQA uses a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this. More details are in the Appendix (Sec. I).
FuSIC-KBQA Patidar et al. (2024) is the SoTA model for few-transfer for KBQA with only answerable questions. It comprehensively outperforms a host of supervised and LLM-equipped KBQA models, such as TIARA Shu et al. (2022), Pangu Gu et al. (2023), KB-Binder Li et al. (2023), gf-LLM Shu and Yu (2024), adapted for few-shot transfer. The logical form generation prompt of FuSIC-KBQA is adapted for unanswerability in the instruction and also the few-shots, which now contain unanswerable exemplars. Specifically, we use 5 shots, of which 2 are answerable, 1 is data-related unanswerable, 1 is schema-related unanswerable, and 1 has missing mentioned entity. This adaption also eliminates execution-guided feedback (EGF) from FuSIC-KBQA, which, being based entirely on empty answer check, is unusable for unanswerability. We name this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, where FUn is removed.
Recall that FUn-FuSIC has two mechanisms for addressing unanswerability via LLM prompting. The first is providing few-shots with unanswerability for logical form generation (PUn). The second is using $n$ iterations of FUn. For a fair comparison, all LLM equipped models should have the same LLM “budget” – equivalently, cumulative prompt size. For this reason, we use FUn-FuSIC with zero-shot generation and 4 iterations of FUn, which roughly equals the prompt length of 5-shot FUn-FuSIC-U.
We use gpt-4-0613 as the LLM for both FUn-FuSIC and FuSIC-KBQA-U. The FUn-FuSIC framework allows flexible use of supervised retrievers. For meaningful comparison, we adapt and use RetinaQA as the retriever for both FUn-FuSIC and FuSIC-KBQA-U. More details about the retriever along with details of compute infrastructure are in the Appendix (Sec. H).
@@
Following existing literature Patidar et al. (2023); Faldu et al. (2024), we evaluate answers using regular and lenient F1 (F1(R) and F1(L)). F1(L) additionally gives credit for the original answer for the complete KB.
For logical form evaluation, the existing evaluation measure (EM) Ye et al. (2022) works only for s-expression as the language. Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL and RetinaQA in s-expression, we propose a new measure (EM-s) that works for both.
EM-s considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answer sets upon KB execution. Since this is a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable, and found $>98\%$ agreement.
More details are in the Appendix (Sec. C).
@@
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on WebQSP $\rightarrow$ GraphQAbility. sc denotes replacing self consistency for unanswerability with self consistency. Other rows remove checks for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
@@ 4.2 Unanswerability Setting
We first address research question R1. Performances of the different models for the few-shot transfer setting with unanswerability are recorded in Tab. 1. We make the following observations. (a) For answerable questions, FuSIC-U and FUn-FuSIC significantly outperform RetinaQA. Across the two datasets, FUn-FuSIC mostly performs better than FuSIC-U (both F1 and EM-s). (b) For unanswerable questions, FuSIC-U performs well in terms of the F1 scores, but not EM-s, indicating empty answers with incorrect logical forms. RetinaQA performs well in terms of F1 scores and EM-s. FUn-FuSIC performs slightly worse than RetinaQA for WebQSP $\rightarrow$ GrailQAbility, but significantly better for WebQSP $\rightarrow$ GraphQAbility. (c) Overall, FUn-FuSIC significantly outperforms both baselines in terms of EM-s, and is slightly below FuSIC-U in terms of F1(L) for WebQSP $\rightarrow$ GrailQAbility. This establishes usefulness of FUn for few-shot transfer KBQA with unanswerability.
We next briefly address research question R2.
Performance of different models for different categories of unanswerability are recorded in Tab. 2. While all models struggle with data-level unanswerability (EM-s), RetinaQA performs the worst and FuSIC-U is the best. For schema-level unanswerability, FUn-FuSIC is the best in one dataset, while RetinaQA is the best in the other. This indicates that more work is needed for robustness across unanswerability categories, and simultaneous use of PUn (prompting) and FUn (feedback) is the most promising approach.
We next address research question R3. Tab. 3 records performance for different ablations of FUn-FuSIC. The biggest benefit, for both answerable and unanswerable questions, comes from question logical form disagreement feedback. Feedback on KB inconsistency and answer incompatibility also make significant contributions to the performance.
@@
Table 4: Performance using F1 of different models for few-shot KBQA transfer with only answerable questions. The models in the first two rows assume knowledge of answerability of all questions, while the those in the next two do not.
@@ 4.3 Answerable Setting
We now address answerable-only KBQA and research question R4. We use two of the datasets used in existing literature for this setting, including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop) Patidar et al. (2024) [^3] https://github.com/dair-iitd/FuSIC-KBQA/. For consistency with existing literature, here we use TIARA Shu et al. (2022) as the retriever for all models.
This setting can be evaluated in two ways. In the first setting, the models are aware that all questions are answerable. This is the setting studied in the KBQA literature, and FuSIC-KBQA is the established SoTA model outperforming a host of supervised and LLM-based models adapted for the task Patidar et al. (2024). We appropriately adapt FUn-FuSIC as discussed earlier in Sec. 3.2.
The first two rows in Tab. 4 record performance for this setting. We see that FUn-FuSIC with self-consistency significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This establishes the usefulness of the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks.
In the more practical second setting, which has not been studied before, although all questions are answerable, the models are not aware of this and make predictions assuming unanswerability. For this we evaluate FuSIC-U and FUn-FuSIC as in Sec. 4.2. The bottom two rows of Tab. 4 record their performance. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability do not exist.
@@ 4.4 Error Analysis
We have performed detailed error analysis for FUn-FuSIC. For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions with EM-s $<1$ (46.2% of questions). There are two high-level error categories. (E1) Retrieval Error: At least one necessary KB element (class, relation, entity) is missing in the retrieved KB context (23.4% of questions). (E2) Generation Error: Despite correct retrieval, the generated logical form is incorrect (22.8% of questions). Generation errors can be further sub-categorized. (E2a) The gold logical form is NK, but some (non-NK) logical form is predicted (8.4% of questions). (E2b) A (non-NK) gold logical form exists, but NK is predicted, (4.6% of questions). (E2c) The gold and the predicted logical forms are non-NK, but different (9.8% of questions). We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms in a complete KB. Incompleteness eliminates one of these, so that the gold logical form is NK, while the model identifies the other still existing non-NK logical form. (2) FUn is unable to generate the gold logical form within its maximum number of iterations ($n=4$). (3) FUn generates the gold logical form but it is rejected incorrectly by the question logical form disagreement checker.

### v2

@@ 4 Proposed Approach: FUn-FuSIC
Unless some logical form passes all checks and is therefore returned (Algo. 2 line 6), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so. For our specific suite of weak verifiers, $l^{(i)}$ is added to $L$ if it passes one of V4b ($A$ is non-empty) as for LF3 for KB1, or V3 ($l^{i}$ is equivalent to $q$), as for LF3 for KB2.
@@
If FUn fails to generate a single confident logical form at the end of $n$ iterations, as for KB1 and KB2, we use scUn to decide if the best candidate $l^{*}\in L$ has sufficient confidence. If scUn detects sufficient confidence, as for KB2, it outputs ($l=l^{*}$, $A=A^{*}$), where $A^{*}$ is the answer obtained upon executing $l^{*}$ and may be NA. If on the other hand, scUn decides insufficient confidence, as for KB1, it outputs ($l=$ NK, $A=$ NA). For identifying the consensus choice from $L$, one possibility is self-consistency Wang et al. (2023); Chen et al. (2023a) i.e., obtaining the answer for each $l\in L$, and returning those with the most common answer. Self-consistency requires some answer to accumulate enough probability by aggregation over reasoning paths. However, for unanswerable questions, no single answer accumulates sufficient probability, and self-consistency returns some low probability answer.
To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$). If so, scUn uses LLM prompting to select the most appropriate supporting logical form $l^{*}\in L$ considering $q$, and outputs ($l=l^{*}$, $A=A^{*}$).
However, for KB1 in the example, the 3 logical forms among the candidates have 3 different answers, and therefore ($\lfloor\frac{|L|}{2}\rfloor=1$) no consensus emerges for non-empty answers. Then scUn considers logical forms from $L$ that agree on $A=$ NA. If there are multiple such candidates, scUn selects the most suitable candidate $l^{*}$, again using LLM prompting, and outputs ($l=l^{*}$, $A=$ NA). If there is no such candidate, scUn outputs ($l=$ NK, $A=$ NA). For KB2 in the example, scUn selects LF3 – the only logical form with empty answer. Further details on scUn are in the Appendix (Sec. A.14). .
@@ 5 Experiments
Pangu 54.5 43.8 31.23 29.6 83.8 80.4 58.0 53.4 33.0 30.3 26.4 76.5 74.8 39.6
FuSIC-KBQA-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
KB-Binder 43.7 33.0 19.5 16.5 67.9 66.5 49.5 44.3 36.1 27.5 21.6 61.0 61.0 50.7
FUn-FuSIC 76.6 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
@@
We now present experimental evaluation of FUn-FuSIC. For few-shot KBQA transfer with unanswerability, we address the following research questions. (R1) How does FUn-FuSIC compare against SoTA models for KBQA, suitably adapted for this setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) How do the different components of FUn-FuSIC contribute to its performance? In addition, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
@@ 5.1 Experimental Setup
For few-shot KBQA transfer, available datasets have only answerable questions Patidar et al. (2024), while our task needs the target dataset to contain unanswerable questions as well. So we construct our own transfer datasets, augmenting existing ones. GrailQAbility is the only available KBQA dataset with unanswerable questions Patidar et al. (2023). This was carefully designed starting from GrailQA Gu et al. (2021), which has only answerable questions, by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries. We use this as one of our targets.
GraphQA Su et al. (2016) is another popular KBQA dataset. This has the same back-end KB (Freebase) as GrailQA. We create our second target dataset using GraphQA, by replacing its KB with the modified KB in GrailQAbility. This introduces unanswerability into GraphQA questions and we label these appropriately as schema-level or data-level unanswerable. We name this dataset GraphQAbility. We create the two test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from their test sets. We create few shots by selecting 100 questions (50 answerable and 50 unanswerable) uniformly at random from the dev set and train set respectively.
@@
WebQSP training set has 2,858 real user questions, which are manually annotated with logical forms. This is quite different from GraphQA and GrailQA which contain algorithmically generated logical forms, verbalized by crowd-workers. Because of this difference in nature of questions, these are challenging transfer datasets, beyond the unaswerability gap.
@@
Table 2: Model performance for categories of unanswerable questions. FuSIC-U is short hand for FuSIC-KBQA-U.
@@
As few-shot transfer for KBQA with unanswerability is a novel task, there are no existing baselines. For in-domain KBQA with unanswerability, RetinaQA Faldu et al. (2024) and Pangu Gu et al. (2023). which has been adapted for unanswerable questions, are the two SoTA models. For these, we use the available code. [^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in Appendix A.10.
FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions. Instead of retrieve-then-generate, KB-Binder Li et al. (2023) follows a generate-then-ground approach. It is the SoTA for in-domain few-shot KBQA, and overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al. (2024). We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA since source code is not available. To adapt these two baselines for unanswerability, for fair comparison, we modify their logical form generation prompt in the same fashion as PUn for FUn-FuSIC. Additionally, for FuSIC-KBQA, we remove execution-guided feedback (EGF) since it fails for unanswerability. We denote this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, without FUn. More details about use of KB-Binder are in the Appendix A.12.
We use $\mathcal{L}=$gpt-4-0613 for all three models. For a fair comparison, we provide all three LLM-equipped models the same aggregated prompt limit for a question. To achieve this, to FUn-FuSIC we provide zero-shot generation and $n=4$ FUn iterations, to FuSIC-KBQA-U 5-shot generation, and to KB-Binder 25-shot generation.
Though FUn-FuSIC and FuSIC-KBQA allow flexible use of multiple supervised retrievers, for meaningful comparison with RetinaQA, we adapt RetinaQA as retriever for FUn-FuSIC and FuSIC-KBQA-U. More details about KB-Binder and FuSIC-KBQA’s generation is in Appendix A.12, and details of FuSIC-KBQA’s retriever along compute infrastructure are in Sec. A.9.
@@
Since our primary task is generation of logical forms, our primary focus in evaluation is on logical forms as well. For logical form evaluation, the existing exact match measure (EM) Ye et al. (2022) works only for s-expression as the language. Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression, we propose a new measure that works for both languages. This measure, EM-s, considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answers upon execution. This is a necessary but not sufficient condition for logical form equivalence. So, we compared EM-s with EM in settings where both are applicable, and found $>98\%$ agreement. More details are in the Appendix A.3.
Since logical form evaluation is not completely accurate (logical forms different from the gold one may still be correct), following existing literature Patidar et al. (2023); Faldu et al. (2024), we also evaluate answers using regular F1 (F1(R)), also lenient F1 (F1(L)), which does not penalize the original answer for the complete KB.
@@
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on subset of WebQSP $\rightarrow$ GraphQAbility. scUn $\Rightarrow$ sc denotes replacing scUn with self consistency. Other rows remove verifiers for syntax error (w/o syntax) (V1), KB inconsistency (w/o kb-inc) (V2), question logical form disagreement (w/o q-lf) (V3) and answer incompatibility (w/o ans-inc) (V4).
@@ 5.2 Unanswerability Setting
We first address research question R1. Performances of the different models for the few-shot transfer setting with unanswerability are recorded in Tab. 1. Note that the Overall columns determine superiority of one model over another, while the Answerable and Unanswerable columns provide further drill-down for analysis.
First, we observe that FUn-FuSIC significantly outperforms all baselines in terms of EM-s, and performs at par with FuSIC-KBQA-U and significantly better than all other models in F1. However, the other LLM-equipped models do not far surpass supervised models. In fact, all 4 models perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility. This establishes usefulness of FUn+scUN for few-shot transfer KBQA with unanswerability. Secondly, each model trades off performance differently between answerable and unanswerable questions. RetinaQA, Pangu and also KB-Binder fare better for unanswerable questions, while FUn-FuSIC and FuSIC-KBQA-U fare better for answerable ones. However, FUn-FuSIC achieves the best balance between the two.
We next briefly address research question R2. Performance of different models for different categories of unanswerability are recorded in Tab. 2. Models struggle to fare well across both data-level and schema-level unanswerability. FuSIC-KBQA-U performs the best for data-level while performing poorly (in terms of EM-s) for schema-level. Conversely, RetinaQA performs well for schema-level, but has poor data-level EM-s. FUn-FuSIC performs poorer for data-level compared to schema-level but achieves the best balance by far between the two categories among all models.
We next address research question R3. Tab. 3 records performance for different ablations of FUn-FuSIC. Note that this experiment is performed on a subset of the test data (50 questions drawn randomly from the each of the answerable and unanswerable categories). The biggest benefit, for both answerable and unanswerable questions, comes from Question - Logical Form Disagreement verifier. KB Inconsistency and Answer Incompatibility verifiers also make significant contributions to the performance. This demonstrates the usefulness of weak verifiers. Replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
@@
Table 4: Performance using F1 of different models for few-shot KBQA transfer with only answerable questions. The models in the top block have prior knowledge of answerability, while those in the bottom block do not.
@@ 5.3 Answerable Setting
We now address research question R4 (answerable-only KBQA). We use two datasets from existing literature Patidar et al. (2024), including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop). [^5] https://github.com/dair-iitd/FuSIC-KBQA/ For consistency with existing literature, here all models use TIARA Shu et al. (2022) as the retriever.
This setting admits two sub-settings: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this prior knowledge.
Setting (A) has been studied for KBQA Patidar et al. (2024), and FuSIC-KBQA is the established SoTA model, outperforming a host of supervised and LLM-based models adapted for the task. In this setting, FUn-FuSIC requires three simple modifications. (i) PUn is replaced with prompt for answerability, (ii) In FUn, category of V4b (empty answer) changes from weak verifier to strong verifier, and (iii) scUn is replaced by self-consistency.
The first two rows in Tab. 4 record performance for setting (A). FUn-FuSIC significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for KBQA transfer, even without unanswerability.
In the realistic second setting (B) – not studied before – the models make predictions assuming unanswerability. Here, we evaluate FuSIC-KBQA-U and FUn-FuSIC as in Sec. 5.2. The bottom two rows of Tab. 4 record their performance. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability are not available.
@@ 5.4 Error Analysis
EM-s $<1$ 46.2
Retr. Err. 23.4
Gen. Err. 22.8
$l^{*}=$ NK, ${\hat{l}\neq}$ NK 8.4
$l^{*}\neq$ NK, ${\hat{l}=}$ NK 4.6
$l^{*}\neq$ NK, ${\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
Table 5: FUn error analysis on WebQSP $\rightarrow$ GraphQAbility. $l^{*}$ ${\hat{l}}$ denote gold & generated logical forms. Retrieval error means retrieval $r$ is missing $\geq 1$ KB elements (class, relation, entity) necessary for $l^{*}$. Generation error implies ${\hat{l}}\neq l^{*}$ despite correct retrieval.
For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions whose logical forms are judged incorrect (EM-s $<1$). See results in Tab. 5. We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms $l_{1}$ and $l_{2}$ in the original complete KB, though only one is recognized as the gold ($l^{*}=l_{1}$). Deletion to introduce unanswerability eliminates $l_{1}$, so we update $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized. (2) $l^{*}=l_{1}$, ${\hat{l}}=l_{2}$, where $l_{1}\neq l_{2}$ while being semantically equivalent, but are incorrectly judged non-equivalent by EM-s. (3) FUn is genuinely unable to generate $l^{*}$ or any semantic equivalent within its iteration limit.
We also estimated the accuracy of specific weak verifiers by manual analysis of $\sim$100 instances from test. Verifier V3 for Question - Logical Form Disagreement has accuracy of 93.6% overall, with its Query to Natural Language Back-translation component having 90%. Verifier V4b for Empty Answer has 67.8% (this is expected since $\sim$25of questions have ($l^{*}\neq$ NK, $A^{*}=$ NA).